The following information is for the Wetherby Company.

2016 2015

Units manufactured 60,000 54,000

Units of materials used 144,000 124,000

Number of labor hours used 200,000 180,000

Cost of materials per unit $40 $38

Direct labor wage rate per hour $50 $44

1. Compute the partial operational productivity measures for 2015 and 2016

. 2. Compute the partial financial productivity ratios for 2015 and 2016.

3. Separate the changes of the partial financial productivity ratios from 2015 to 2016 into productivity change, input price change, and output change.
Answer #1

Answer 1

For 2015

Partial Operating productivity measures ( Material) = Units Manufactured/Units of material used

= 54000/124000

= 0.4355

Partial Operating productivity measures ( Labour) = Units Manufactured/No of labour hours

= 54000/180000

= 0.30

For 2016

Partial Operating productivity measures ( Material) = Units Manufactured/Units of material used

= 60000/144000

= 0.4167

Partial Operating productivity measures ( Labour) = Units Manufactured/No of labour hours

= 60000/200000

= 0.30

Answer 2

For 2015

Partial Financial productivity ratios ( Material) = Units Manufactured/Amount spent on materials used

= 54000/(124000*38)

= 0.0115

Partial Financial productivity ratios ( Labour) = Units Manufactured/Amount spent on labour

= 54000/(180000*44)

= 0.0068

For 2016

Partial Operating productivity measures ( Material) = Units Manufactured/Amount spent on materials used

= 60000/(144000*40)

= 0.0104

Partial Operating productivity measures ( Labour) = Units Manufactured/Amount spent on labour

= 60000/(200000*50)

= 0.0006